# Consistency Flow Model Achieves One-step Denoising Error Correction Codes

### v1

@@ 4.1 Error Correction Consistency Flow Property
However, in ECC domain, Binary Cross Entropy (BCE) is the commonly used divergence measure. By Proposition 4.1, we demonstrate that our formulation in Eq. 8 serves as an upper bound on the vanilla consistency loss in terms of Total variation distance, directly optimizing for the Error Correction Consistency property.
@@
Let $\mathcal{L}_{\text{Standard-CM}}$ be defined by the Total Variation distance, $\text{TV}(\cdot,\cdot)$, and $\mathcal{L}_{\text{EC-CM}}$ be defined by the Binary Cross Entropy, $\text{BCE}(\cdot,\cdot)$. For any timesteps $t,r$ and ground truth $\mathbf{x}_{0}$, the following semi–triangle inequality holds:
@@
This fundamental change requires all decoding trajectories mapping to their origin codeword, as demonstrated in Figure 1, and the learned solution distribution is expected to center on $\mathbf{x}_{0}$.
4.2 Differential Consistency Condition for Smooth Trajectory
@@
The training objective of a consistency model is to learn a function $f_{\theta}(\mathbf{x}_{t},t)$ that maps any point on a trajectory back to its origin $\mathbf{x}_{0}$. This is enforced by both satisfying the self-consistency property and the boundary condition, which is rooted in the differential equation $\frac{\mathrm{d}f}{\mathrm{d}t}=0$. Following Eq. 10 and 11 in [geng2024consistency], the consistency function $f_{\theta}$ is parameterized to satisfy these conditions:
@@
A critical problem arises when applying the consistency framework to ECC decoding tasks: the time variables $t$ and $r$, which represent noise levels, are not directly observable from the received signals. A seemingly natural solution, proposed in DDECC [choukroundenoising], is to use the sum of syndrome error, $e_{t}$, as a measurement of the noise level, since an error count of zero ($e_{t}=0$) indicates a valid codeword. The hard syndrome is computed as $s=Hy_{b}^{\top}$, where $H$ is the parity-check matrix. The sum of syndrome error, $e_{t}=\sum_{i}s(\mathbf{x}_{t})_{i}$, is then the sum of binary syndrome bits.
@@
However, naively replacing the continuous time variables $t$ and $r$ in Eq. 11 with their discrete, integer-valued counterparts $e_{t}$ and $e_{r}$ violates the core assumption of a smooth, differential trajectory. As we demonstrate in Figure 3, the trajectory defined by the syndrome error is highly non-smooth and clustered: A small change in the noisy signal can cause an abrupt jump in the syndrome error count, invalidating the finite-difference approximation; And it is common for two different noisy signals, $\mathbf{x}_{t}$ and $\mathbf{x}_{r}$, to have the same syndrome error count ($e_{t}=e_{r}$), leading to instability during training.
@@
Therefore, the discrete and non-smooth nature of the syndrome error makes it an unsuitable conditioning variable for standard consistency model training in ECC, which necessitates a smooth and differential measure of noise level along the denoising trajectory. We then propose replacing this discrete error sum condition with a continuous and differentiable alternative. Inspired by [lau2025interplaybeliefpropagationtransformer], we introduce the soft syndrome as the basis for the consistency noise condition. The soft syndrome, $s^{\dagger}$, is a fully differentiable function that leverages the log-likelihood ratios of the received signal $\mathbf{x}_{t}$: and offers a continuous measure of how close each parity-check equation is to being fulfilled.
Algorithm 1 Error Correction Consistency Training
@@
Model $f_{\bm{\theta}}$, parity-check matrix $\mathbf{H}$, learning rate $\eta$, syndrome weight $\lambda$, denoising steps NN, time scaling factor $\alpha$.
for training batch $\mathbf{x}_{0}$ do
@@
  $t\sim\mathcal{U}\{1,\dots,N\},\,r=\alpha t$ $\triangleright$ Sample forward diffusion steps
  $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ $\triangleright$ Sample Gaussian noise
@@
  $\mathbf{x}_{t}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{t}}\cdot\bm{\epsilon},\mathbf{x}_{r}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{r}}\cdot\bm{\epsilon}$ $\triangleright$ Generate noisy signals via forward process
@@
  $e^{\dagger}_{t}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H),\,\,e^{\dagger}_{r}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{r},H)$ $\triangleright$ Calculate soft-syndrome
  $\mathcal{L}_{\text{EC-CM}}\leftarrow\mathcal{L}_{\text{BCE}}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),\mathbf{x}_{\text{0}})+\mathcal{L}_{\text{BCE}}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),\mathbf{x}_{\text{0}})$
@@
  $\mathcal{L}_{\text{Total}}\leftarrow\mathcal{L}_{\text{Consistency}}+\lambda\cdot\left(\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),H)+\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),H)\right)$
  $\bm{\theta}\leftarrow\bm{\theta}-\eta\cdot\nabla_{\bm{\theta}}\mathcal{L}_{\text{Total}}$
end for
@@
$\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H)$ $=-e^{\dagger}_{t},$ (13)
This soft-syndrome error condition is computed as the binary cross-entropy between the estimated syndrome and the all-zero syndrome, which requires valid codewords to satisfy all parity-check equations. We use the mean-field approximation to provide differential conditions to estimate the probability of satisfying zero-syndrome conditions in Eq. 12:
$s_{j}^{\dagger}=\frac{1}{2}-\frac{1}{2}\prod_{\{i\mathrel{\mathop{\ordinarycolon}}H_{j,i}=1\}}\Bigl(2\cdot\mathrm{sigmoid}\bigl(\frac{2\mathbf{x}_{t,i}}{\sigma^{2}}\bigr)-1\Bigr),\quad\Pr(S_{j}^{\dagger}=0)=1-s_{j}^{\dagger},$ (14)
where $x_{t,i}$ is the $i$-th position of the received codeword $x_{t}$, the noise level $\sigma$ is given to the decoder at the receiver’s side. The soft-syndrome is zero if and only if the codeword is valid, yet it varies smoothly and continuously with the received signal $\mathbf{x}_{t}$. By using $e^{\dagger}_{t}$ as the time condition, we provide the consistency model with a smooth, differentiable trajectory from a noisy signal to a valid codeword, resolving the instability and degeneracy issues of the hard error sum and enabling stable training.
4.3 Error Correction Consistency Flow Training Dynamics
@@
where $d(\cdot,\cdot)$ is a distance metric, such as Binary Cross-Entropy (BCE) in this work, and $\lambda$ is a hyperparameter that weights the syndrome regularization term. Once trained according to Algorithm 1, the learned consistency function $f_{\theta}$ can decode the noisy received signal $y$ in a one step: $\hat{\mathbf{x}_{0}}=f_{\theta}(y,e^{\dagger}_{t})$, as shown in Appendix 2, Algorithm 2.
@@ 5 Numerical Results
Evaluation Metrics. We evaluate decoding performance using two standard metrics following established benchmarks [choukroun2022error, choukroundenoising, parkcrossmpt]: Bit Error Rate (BER) and Frame Error Rate (FER). BER measures the fraction of individual bits that are incorrectly decoded. FER (also known as Block Error Rate, BLER) measures the fraction of entire codewords that contain one or more bit errors. Concerning the latency factor, we evaluate computational efficiency by reporting inference time and throughput (decoded samples per second).
Baselines. We numerically compared the results with multiple baselines for the decoding task, including: 1) Conventional BP-based decoders: BP [bennatan2018deep] and ARBP [nachmani2021autoregressive]. 2) Auto-regressive model-free decoders: ECCT [choukroun2022error], and CrossMPT [parkcrossmpt]. 3) Denoising diffusion model-free decoders: DDECC [choukroundenoising].
@@
Overall Performance. Following established benchmarks [choukroun2022error, choukroundenoising, parkcrossmpt], we conducted a decoding performance comparison measured in $\ln(\text{BER})$. Our method was evaluated versus two classes of decoders: conventional BP-based algorithms (BP and ARBP) and model-free neural decoders (ECCT, CrossMPT, and DDECC). To ensure a fair comparison, all neural models were implemented with a fixed architecture ($N=6$ layers, $d=128$ hidden dimensions). Furthermore, to ensure statistical significance, each simulation was run until at least 500 error codes were observed, under a maximum of $10^{8}$ test instances.
As shown in Table 1, our proposed ECCFM framework consistently achieves better performance across the tested code families, including BCH, Polar, LDPC, CCSDS and MacKay with different code rates $(n,k)$. In the test cases, ECCFM achieves the best or second-best BER, performing better than autoregressive and diffusion-based neural decoders and showing considerable gain over POLAR codes. This result indicates the applicability and effectiveness of the proposed consistency-based training approach.
Performance on Longer Codes. To further evaluate scalability, we conducted a focused evaluation on longer codes commonly used in practical communication systems: LDPC(n=204,k=102)(n=204,k=102), LDPC(n=529,k=440)(n=529,k=440), WRAN(n=384,k=320)(n=384,k=320), and Polar(n=512,k=384)(n=512,k=384). All methods were implemented with the same model architecture ($N=6,d=128$). As illustrated in Figure 4, ECCFM improves upon the scores of the neural net baselines across a range of SNRs (2 dB to 6 dB), highlighting the scalability and robustness of ECCFM for those challenging decoding tasks. Additional results on other high-length codes are presented in Appendix 7.5.7, which also suggest inference speed gain compared to the baselines.
Performance on Rayleigh Fading Channel. While our proposed ECCFM training dynamics and the definition of soft-syndrome are intrinsic to the code structure and independent of the underlying channel model, we validate our approach on Rayleigh fading channels to demonstrate robustness. We follow the same experimental settings described in [choukroun2022error, park2024crossmpt]. In contrast to the AWGN channel, the received signal is defined by: $y=hx+z$, where $h$ is an $n$-dimensional i.i.d. Rayleigh distributed vector with a scale parameter $\alpha=1$ and $z\sim N(0,\sigma^{2})$. In Figure 6, our method achieved competitive performance compared to baselines across different SNRs on Polar and LDPC. These numerical results confirm that our proposed ECCFM framework generalizes effectively to non-Gaussian channels without fundamental modifications.
@@
Inference Time and Throughput Comparison. A key benefit of ECCFM is its ability to perform high-fidelity decoding in only one step. To quantify this efficiency gain, we measured inference time (total seconds to decode $10^{5}$ samples) and throughput (samples decoded per second). As shown in Figure 5, ECCFM demonstrates a speed advantage over diffusion-based methods such as DDECC, achieving speedups of over 30x for short codes and 100x for medium-to-long codes. This disparity arises because diffusion models require several iterative denoising steps for inference, a computational cost that scales with code complexity as detailed in Appendix 7.5.8. Notably, ECCFM achieves decoding speeds comparable to the fastest auto-regressive baseline (CrossMPT) due to its one-step nature. Therefore, ECCFM matches the competitive performance of denoising diffusion decoders while operating at the high throughput of single-step auto-regressive decoders.
Ablation Study: Model-Agnostic Property of ECCFM. We discussed that ECCFM is a model-agnostic training framework, i.e., its performance could be preserved over different neural network architectures. To test this, we conducted an ablation study where we decoupled our framework from the cross-attention transformer backbone conducted before. Specifically, we took the underlying architecture of the ECCT baseline and trained it using our proposed ECCFM training objective. We then compared this model directly against the original ECCT, which uses the same architecture. The results in Table 2 show that applying the ECCFM training objective yields improvement in $\ln(\text{BER})$ over the standard ECCT, with comparable performance versus the iterative denoising DDECC method. These results demonstrate that ECCFM’s property of model-agnostic and its benefits are attributable to the consistency-based training framework itself.
Table 2: Performance comparison of ECCFM versus the standard ECCT and DDECC, using an identical ECCT backbone on POLAR and LDPC codes.
Architecture ECCT Backbone
Code Type Parameters ECCT DDECC ECCFM(ECCT)
4 5 6 4 5 6 4 5 6
POLAR (64,32) 6.87 9.21 12.15 7.04 9.44 12.70 7.12 9.77 12.71
(64,48) 6.21 8.31 10.85 5.93 8.00 10.44 6.38 8.55 11.23
(128,64) 5.79 8.45 11.10 7.71 11.40 13.85 7.32 11.03 14.87
(128,86) 6.29 8.98 12.82 7.61 10.50 13.88 7.18 10.17 15.02
(128,96) 6.30 9.04 12.40 7.14 10.31 13.66 6.86 9.94 13.83
LDPC (121,60) 5.12 8.21 12.80 5.42 9.11 13.82 5.55 8.86 13.97
(121,70) 6.30 10.11 15.50 6.91 11.02 17.15 6.87 11.21 16.13
(121,80) 7.27 11.21 17.02 7.61 11.89 16.18 7.80 12.03 17.95
Ablation Study: Necessity of Soft-syndrome Time Condition in Consistency Training. We further investigated the necessity of using soft-syndrome as a time condition for building the smooth trajectory. We trained ECCFM with either a hard-syndrome time condition or a soft-syndrome time condition. In Appendix 7.5.5, Table 7, our experiments revealed that using soft-syndrome as time condition is necessary for successful training of ECCFM. The continuous nature of the soft-syndrome provides a smooth reverse trajectory to learn, whereas models conditioned on the hard syndrome failed to converge.

### v2

@@ 4.1 Syndrome-Flow Consistency Property
In the ECC domain, we use Binary Cross Entropy (BCE) for the distance measure $d(\cdot,\cdot)$. Our Proposition 4.1 theoretically demonstrates that our optimization objective, $\mathcal{L}_{\text{EC-CM}}$, also serves as an upper bound on the standard consistency loss measured by Total Variation distance. This confirms that our more direct approach inherently enforces self-consistency.
@@
Let $\mathcal{L}_{\text{Standard-CM}}$ be defined by the Total Variation distance, $\text{TV}(\cdot,\cdot)$, and $\mathcal{L}_{\text{EC-CM}}$ be defined by the Binary Cross Entropy, $\text{BCE}(\cdot,\cdot)$. For any timesteps $t,r$ and ground truth $\mathbf{x}0$, the following semi–triangle inequality holds:
@@
This change in the training objective compels all decoding trajectories to map directly to their origin This fundamental change requires all decoding trajectories mapping to their origin codeword, as demonstrated in Figure 1, and the learned solution distribution is expected to center on $\mathbf{x}_{0}$.
4.2 Soft-syndrome Condition for Smooth Trajectory
@@
The training objective of a consistency model is rooted in the differential equation $\frac{\mathrm{d}f}{\mathrm{d}t}=0$. Following Eq. 10 and 11 in [12], the consistency function $f_{\theta}$ is parameterized to satisfy these conditions:
@@
A critical problem arises when applying the consistency framework to decoding tasks: the time variables $t$ and $r$, which measure noise levels, are not directly observable from the received signals. A natural solution, proposed in DDECC [6], is to use the sum of syndrome error, $e_{t}$, as a measurement of the noise level, since $e_{t}=0$ indicates a valid codeword. The hard syndrome is computed as $s=Hy_{b}^{\top}$, where $H$ is the parity-check matrix. The sum of syndrome error, $e_{t}=\sum_{i}s(\mathbf{x}_{t})_{i}$, is then the sum of binary syndrome bits.
@@
However, replacing the continuous time variables $t$ and $r$ in Eq. 6 with their discrete, integer-valued counterparts $e_{t}$ and $e_{r}$ violates the core assumption of a smooth, differential trajectory in consistency training. As we demonstrate in Figure 3, the trajectory defined by the syndrome error is highly non-smooth and clustered. A small change in the noisy signal can cause an abrupt jump in the syndrome error count, invalidating the finite-difference approximation and leading to instability during training in 3.
@@
Therefore, the discrete and non-smooth nature of the syndrome error makes it an unsuitable conditioning variable for consistency training in ECC, which necessitates a smooth and differential measure of noise level along the denoising trajectory. We then propose replacing this discrete error sum condition with a continuous and differentiable alternative. Inspired by [20], we introduce the soft syndrome as the basis for the consistency noise condition. The soft syndrome, $s^{\dagger}$, is a fully differentiable function that leverages the log-likelihood ratios of the received signal $\mathbf{x}_{t}$: and offers a continuous measure of how close each parity-check equation is to being fulfilled.
Algorithm 1 Syndrome-Flow Consistency Training
@@
1:Model $f_{\bm{\theta}}$, parity-check matrix $\mathbf{H}$, learning rate $\eta$, syndrome weight $\lambda$, denoising steps NN, time scaling factor $\alpha$.
2:for training batch $\mathbf{x}_{0}$ do
@@
3:  $t\sim\mathcal{U}\{1,\dots,N\},\,r=\alpha t$ $\triangleright$ Sample forward diffusion steps
4:  $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ $\triangleright$ Sample Gaussian noise
@@
5:  $\mathbf{x}_{t}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{t}}\cdot\bm{\epsilon},\mathbf{x}_{r}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{r}}\cdot\bm{\epsilon}$ $\triangleright$ Generate noisy signals via forward process
@@
6:  $e^{\dagger}_{t}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H),\,\,e^{\dagger}_{r}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{r},H)$ $\triangleright$ Calculate soft-syndrome
7:  $\mathcal{L}_{\text{EC-CM}}\leftarrow\text{BCE}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),\mathbf{x}_{\text{0}})+\text{BCE}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),\mathbf{x}_{\text{0}})$
@@
8:  $\mathcal{L}_{\text{Total}}\leftarrow\mathcal{L}_{\text{Consistency}}+\lambda\cdot\left(\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),H)+\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),H)\right)$
9:  $\bm{\theta}\leftarrow\bm{\theta}-\eta\cdot\nabla_{\bm{\theta}}\mathcal{L}_{\text{Total}}$
10:end for
@@
This soft-syndrome error condition is computed as the binary cross-entropy between the estimated syndrome and the all-zero syndrome. We use the mean-field approximation to provide differential conditions to estimate the probability of satisfying zero-syndrome conditions in Eq. 7:
$s_{j}^{\dagger}=\frac{1}{2}-\frac{1}{2}$ $\prod_{\{i\mathrel{\mathop{\ordinarycolon}}H_{j,i}=1\}}\Bigl(2\cdot\mathrm{sigmoid}\bigl(\frac{2\mathbf{x}_{t,i}}{\sigma^{2}}\bigr)-1\Bigr),$ (8)
$\Pr(S_{j}^{\dagger}=0)=1-s_{j}^{\dagger},$
where $x_{t,i}$ is the $i$-th position of the received codeword $x_{t}$, the noise level $\sigma$ is given to the decoder at the receiver’s side. The soft-syndrome is zero if and only if the codeword is valid, yet it varies smoothly and continuously with the received signal $\mathbf{x}_{t}$. By using $e^{\dagger}_{t}$ as the time condition, we provide the consistency model with a smooth trajectory from a noisy signal to a valid codeword, resolving the instability and degeneracy issues of the hard error sum and enabling stable training.
4.3 Syndrome-Flow Consistency Training Dynamics
@@
where $\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H)=-e^{\dagger}_{t}$, $d(\cdot,\cdot)$ is a distance metric, such as Binary Cross-Entropy (BCE) in this work, and $\lambda$ is a hyperparameter that weights the syndrome regularization term. Once trained according to Algorithm 1, the learned consistency function $f_{\theta}$ can decode the noisy received signal $y$ in a one step: $\hat{\mathbf{x}}_{0}=f_{\theta}(y,e^{\dagger}_{t})$, in Appendix 2, Algorithm 2.
@@ 5 Numerical Results
Evaluation Metrics. We evaluate decoding performance using two standard metrics following established benchmarks [5, 6, 31]: Bit Error Rate (BER) and Frame Error Rate (FER). BER measures the fraction of individual bits that are incorrectly decoded. FER measures the fraction of entire codewords that contain one or more bit errors. Concerning the latency, we evaluate computational efficiency by reporting inference time and throughput (decoded samples per second).
Baselines. We numerically compared the results with multiple baselines for the decoding task, including: 1) Conventional BP-based decoders: BP [1] and ARBP [29]. 2) Transformer-based decoders: ECCT [5], and CrossMPT [31]. 3) Denoising diffusion decoders: DDECC [6].
@@
Overall Performance. Following established benchmarks [5, 6, 31], we conducted a decoding performance comparison measured in $\ln(\text{BER})$. Our method was evaluated versus two classes of decoders: conventional BP-based algorithms (BP and ARBP) and model-free neural decoders (ECCT, CrossMPT, and DDECC). To ensure a fair comparison, all neural models were implemented with a fixed architecture ($N=6$ layers, $d=128$ hidden dimensions). Furthermore, to ensure statistical significance, each simulation was run until at least 500 error codes were observed, under a maximum of $10^{8}$ test instances. As shown in Table 1, our proposed ECCFM framework consistently achieves the best or second-best BER across the tested code families, including BCH, Polar, LDPC, CCSDS and MacKay with different code rates $(n,k)$, and showing considerable gain over POLAR codes.
@@
Performance on Longer Codes. To further evaluate scalability, we conducted a evaluation on longer codes commonly used in practical communication systems: LDPC(n=204,k=102)(n=204,k=102), LDPC(n=529,k=440)(n=529,k=440), WRAN(n=384,k=320)(n=384,k=320), and Polar(n=512,k=384)(n=512,k=384). All methods were implemented with the same model architecture ($N=6,d=128$). As illustrated in Figure 4, ECCFM improves the performance across a range of SNRs (2 dB to 6 dB). Additional results on other high-length codes are presented in Appendix 7.5.8, which also suggest inference speed gain compared to the baselines.
@@
Performance on Rayleigh Fading Channel. While our proposed ECCFM training dynamics and the definition of soft-syndrome are intrinsic to the code structure and independent of the underlying channel model, we validate our approach on Rayleigh fading channels to demonstrate robustness. We follow the same experimental settings described in [5, 31]. In contrast to the AWGN channel, the received signal is defined by: $y=hx+z$, where $h$ is an $n$-dimensional i.i.d. Rayleigh distributed vector with a scale parameter $\alpha=1$ and $z\sim N(0,\sigma^{2})$. In Figure 6, our method achieved competitive performance compared to baselines across different SNRs on Polar and LDPC without modifications.
@@
Inference Time and Throughput Comparison. To quantify the efficiency gain, we measured inference time (total seconds to decode $10^{5}$ samples) and throughput (samples decoded per second). As shown in Figure 6, ECCFM demonstrates a speed advantage over diffusion-based methods such as DDECC, achieving speedups of over 30x for short codes and 100x for medium-to-long codes. This disparity arises because diffusion models require several iterative denoising steps for inference, a computational cost that scales with code complexity as detailed in Appendix 7.5.9. Therefore, ECCFM matches the competitive performance of denoising diffusion decoders while operating at the high throughput of single-step transformer-based decoders.
Ablation Study: Model-Agnostic Property of ECCFM. We discussed that ECCFM is a model-agnostic training framework, i.e., its performance could be preserved over different neural network architectures. We conducted an ablation study where we decoupled our framework from the backbone. Specifically, we used the ECCT baseline architecture and trained it with our proposed ECCFM and compared this model against the original ECCT and DDECC. The results in Appendix 7.5.4, Table 5 show that applying the ECCFM training objective yields improvement in $\ln(\text{BER})$ over the standard ECCT, with comparable performance versus the iterative denoising DDECC method.
Ablation Study: Necessity of Soft-syndrome Time Condition in Consistency Training. We further investigated the necessity of using soft-syndrome as a time condition for building the smooth trajectory. We trained ECCFM with either a hard-syndrome time condition or a soft-syndrome time condition. In Appendix 7.5.6, Table 7, our experiments revealed that using soft-syndrome as time condition is necessary for successful training of ECCFM. The continuous nature of the soft-syndrome provides a smooth trajectory to learn, whereas models conditioned on the hard syndrome failed to converge.